OpenStudy (jimmiejones153):

5(4x + 2) ÷ 7y From the expression above, provide an example of each of the following: sum, term, product, factor, quotient, and coefficient. If any are not present, write "not present."

OpenStudy (jimmiejones153):

can i get help

OpenStudy (supersmart1001):

can you list out what you think each of the terms are and we can go from there

OpenStudy (supersmart1001):

Let's start off with "sum". I'm assuming you know that sum means addition.. So what part of your equation above is showing addition?

OpenStudy (supersmart1001):

(4x+2)

OpenStudy (supersmart1001):

Correct! Okay next is term. A term is a number or a variable, so what in your equation is a number and/or a variable?

OpenStudy (supersmart1001):

Hint: Basically every number/letter in your equation

OpenStudy (jimmiejones153):

ok

OpenStudy (supersmart1001):

do u know the answer now

OpenStudy (supersmart1001):

u can solve 5(4x-2)+9=2(8x+7) right?

OpenStudy (jimmiejones153):

no

OpenStudy (supersmart1001):

Simplifying 5(4x + -2) + 9 = 2(8x + 7)

OpenStudy (supersmart1001):

Reorder the terms: 5(-2 + 4x) + 9 = 2(8x + 7) (-2 * 5 + 4x * 5) + 9 = 2(8x + 7) (-10 + 20x) + 9 = 2(8x + 7)

OpenStudy (supersmart1001):

-10 + 9 + 20x = 2(8x + 7) Combine like terms: -10 + 9 = -1 -1 + 20x = 2(8x + 7) Reorder the terms: -1 + 20x = 2(7 + 8x) -1 + 20x = (7 * 2 + 8x * 2) -1 + 20x = (14 + 16x) Solving -1 + 20x = 14 + 16x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-16x' to each side of the equation.

OpenStudy (supersmart1001):

-1 + 20x + -16x = 14 + 16x + -16x Combine like terms: 20x + -16x = 4x -1 + 4x = 14 + 16x + -16x Combine like terms: 16x + -16x = 0 -1 + 4x = 14 + 0 -1 + 4x = 14 Add '1' to each side of the equation. -1 + 1 + 4x = 14 + 1 Combine like terms: -1 + 1 = 0 0 + 4x = 14 + 1 4x = 14 + 1 Combine like terms: 14 + 1 = 15 4x = 15 Divide each side by '4'. x = 3.75 Simplifying x = 3.75

OpenStudy (supersmart1001):

do u get it now?

OpenStudy (jimmiejones153):

yes

OpenStudy (supersmart1001):

ok